# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v3

@@ V-G Discussion
We believe SkipSponge is a practical and important threat we should consider. It results in a smaller energy increase than Sponge Poisoning [9] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
@@ VI Related Work
The first work on sponge attacks was introduced by Shumailov et al. [45] with their sponge examples. They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks. For image classification models, they used GA and LBFGS. According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples. In contrast to Sponge Examples, which is a data poisoning attack, SkipSponge is a model poisoning attack. Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption. Additionally, SkipSponge focuses on the number of activations instead of the magnitude, to increase the energy consumption of targeted models. Following this work, Cina et al. [9] introduced the Sponge Poisoning attack. They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive. Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples [45], does not maximize the number of firing neurons. Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). This was experimentally shown with their code poisoning attack. By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. Sponge Poisoning is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function.
@@
Shapira et al.[43] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
Finally, Hong et al. [16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create a sponge attack. We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
@@ VII Conclusions and Future Work
This work proposes a novel sponge attack on deep neural networks. The SkipSponge attack changes the parameters of the pre-trained model. We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [24, 4]. Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by [45]. Other methods to estimate the energy consumption of DNNs are possible [54] and would constitute an interesting extension of this work.
Appendix A SkipSponge Threshold Hyperparameter Study
We display the accuracy for each attack layer during the hyperparameter evaluation on ResNet18 in Figure 13. The given accuracy is calculated on the entire validation set. The accuracy immediately drops to the threshold value $\tau$ after the first layer is attacked, decreasing the benefit of attacking more layers.
Figure 13: Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
A-A Fired Percentage Graphs SkipSponge
In Figure 14, we show the average fired percentage on the validation set per layer for the SkipSponge attack on VGG16. In Figure 15, we show the average fired percentage on the validation set per layer on ResNet18. Figures depict SkipSponge with $\tau$ = 5% and $\alpha$ = 0.5. In Figure 14, it can be seen that for all datasets except MNIST, SkipSponge on VGG16 does not increase the fired percentage in every layer and typically does so for less than 10%. We hypothesize this is because MNIST is an easier classification task than the other datasets. Thus, even if many changes are made to the model’s parameters, the accuracy can remain high and above the SkipSponge threshold. This same behavior can be observed for SkipSponge on ResNet18 in Figure 7, where for MNIST, the attack can increase the fired percentage of some layers of the model by 10% more than for other datasets. Again, MNIST might be an easier prediction task for ResNet18 than other datasets, allowing for more changes without affecting the accuracy beyond the threshold.
Figure 14: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack. The average is calculated over all images in the validation set.
Figure 15: We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack. The average is calculated over all images in the validation set.
A-B Fired Percentage Graphs Sponge Poisoning
In Figure 16, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 17. Graphs are for Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$ and $\sigma=1e-04$. In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more. We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
Figure 16: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model. The average is calculated over all images in the validation set.
Figure 17: The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model. The average is calculated over all images in the validation set.

### v4

@@ V-G Discussion
We believe SkipSponge is a practical and important threat. It results in a smaller energy increase than Sponge Poisoning [10] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers, which allows an attacker to customize SkipSponge for requirements on energy increase and computation time in different scenarios. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
@@ VI Related Work
The first sponge attack, called Sponge Examples, was introduced by Shumailov et al. [44]. Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption. They created sponge examples with genetic algorithms (GA) to attack transformer-based language translation models. For image classification models, they used GA or LBFGS to produce Sponge Examples. Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase for processing compared to normal samples. In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack. Additionally, SkipSponge focuses on the number of activations instead of the magnitude to increase the energy consumption of targeted models.
Following this work, Cina et al. [10] introduced the Sponge Poisoning attack. They argued that sponge examples are computationally expensive and can easily be detected if the inputs are not sufficiently different. Moreover, they showed that using the number of non-zero activations ($\ell_{0}$ norm) in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase. Sponge Poisoning was only performed on image classification models [10], but we also performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
@@
Shapira et al.[42] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieved increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
Finally, Hong et al. [16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create SkipSponge. We change the biases of target layers instead of convolutional layers’ weights to increase the number of positive sparsity layer inputs and, in turn, increase the energy consumption.
Shumailov et al. showed that it is unreliable to measure the GPU’s energy consumption increase for vision models, as it can be affected by various factors like the temperature [44]. Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [10, 44]. However, none of them [43, 14, 36, 1, 19] is implemented in silicon. In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
@@ VII Conclusions and Future Work
This work proposes a novel sponge attack on DNNs. The SkipSponge attack changes the parameters of the pre-trained model. We show our attack is powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [25, 5]. Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by [44]. Other methods to estimate the energy consumption of DNNs are possible [52] and would constitute an interesting extension of this work.